QueenMidnight16:

I need a little help identifying additive inverse and commutative property and stuff like that. Pretty sure i should know this by now lol.

kaleighh533:

im not good at explaining this type of stuff but i found this site maybe it can help you better than i can. (: https://www.ck12.org/book/ck-12-middle-school-math-concepts-grade-8/r14/section/2.13/

FexrlessRxby:

For the additive inverse, it's the same number, but with a different sign. So for example: 6 + (-6) = 0 and vice versa. It also helps when solving for x. For example: x + 20 + (-20) = 5 + (-20) If you could do that problem and tell me when you get the answer.

QueenMidnight16:

thank you!

FexrlessRxby:

Of course!

FexrlessRxby:

Also, if you got your answer, can you please close this post?

Ashtiel:

I'm not good with this stuff so It would probably just be better for me to help you with just giving you a website that'll help you https://www.ck12.org/book/ck-12-middle-school-math-concepts-grade-8/r14/section/2.13/